OpenStudy (anonymous):

find the values of the trigonometric functions of the following quadrantal angled; (1)\(\dfrac5/{2}*pi\) (2).\(\dfrac-9{2}*pi\) (3).\(\dfrac235{2}\)

OpenStudy (anonymous):

sorry for above posting here are the questions \(\dfrac{5}{2}pi\) (2)-.\(\dfrac{9}{2}pi\) (3).\(\dfrac{235}{3}pi\)

OpenStudy (math&ing001):

Before you plug them into any trigonometric function make sure to simplify them first: 5/2 pi = 2pi + 1/2 pi - 9/2 pi = -4pi - 1/2 pi Knowing than sin(x+2pi)=sin(x) and same for all trigonometric functions this should make it easier.

OpenStudy (anonymous):

what is the technique for simplifying first @math&ing001

OpenStudy (anonymous):

@amistre64

OpenStudy (math&ing001):

Here's an example : cos(5/2 pi) = cos(2pi + 1/2 pi) = cos(1/2 pi) = 0

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

idon't understand how we do ..i fill v.tough

OpenStudy (anonymous):

What is the question and where is it??

OpenStudy (anonymous):

Okay I got it.. :)

OpenStudy (anonymous):

Do you know how to convert radians into Degrees??

OpenStudy (anonymous):

@waterineyes yes i know conversion .

OpenStudy (anonymous):

If you can convert it, it will look easy to you, otherwise we can work in radians too.. What you want??

OpenStudy (anonymous):

do in radians ..it would be my repeatation

OpenStudy (anonymous):

Are you getting what @math&ing001 saying in his posts?? I suggest you, first listen to him.. If you still don't get it anywhere, then I am here to help you.. :) Okay??

OpenStudy (math&ing001):

\[\sin(\frac{ 5 }{ 2 }\pi)=\sin(\frac{ 4+1 }{ 2 }\pi)=\sin(\frac{ 4 }{ 2 }\pi+\frac{ 1 }{ 2 }\pi)=\sin(2\pi+\frac{ 1 }{ 2 }\pi)=\sin(\frac{ 1 }{ 2 }\pi)=1\]  I can not explain it any more that this.

OpenStudy (anonymous):

expalin me step by step ...math@ing tell me but i can't understand

OpenStudy (anonymous):

Okay, then I start.. You must know firstly, where cos(x) is maximum, for same x, sin(x) will be minimum.. Now, I am saying as for any odd multiple of pi/2, cos(x) will always be 0, this you got??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x) means theta ..he na

OpenStudy (anonymous):

Odd multiple of pi/2 means : pi/2, 3pi/2, 5pi/2 etc... with pi/2, there must come odd digit.. Yes, theta and x is the same thing..

OpenStudy (anonymous):

For x to be odd multiple of pi/2, cos(x) will be 0 and sin(x) will be 1 or -1.. This I will show you later.. Now, let us come to your question.. Your first part is : 5pi/2..

OpenStudy (anonymous):

yes @ water in eyes

OpenStudy (anonymous):

Before this, you must know this also: Complementary of sin is cos and vice versa.. Complementary of tan is cot and vice versa. Complementary of sec is cosec and vice versa..

OpenStudy (anonymous):

what means of complementary